# Industrial Internet of Things Intelligence Empowering Smart Manufacturing: A Literature Review

### 02 Dec 2023 v1

@@ IV-B1 Motivations
The revolution of IIoT networks is inspired by global trends[44]. Firstly, because of more and more smart equipment and staff accessing the network in the domain of smart manufacturing, and the connection between systems, human, processes, locations and devices getting more distributed and complex, the economic value of the network for enterprises are growing. Secondly, the inherent unpredictability of business that the network quickly corrects the transmission modes to adjust to the new services, processes and models. Thirdly, many manufacturing processes are time sensitive and mission critical, which requires the communication time of network to be more precise and predictable[45]. Moreover, advanced networks are expected to offer the promise of the environmental sustainability.
Technology improvements encourage IIoT network revolution as well. The applications and data of enterprises are leaving the premises and moving to the public/private clouds. Part of the monolithic applications in many cases are being modularized into microservices that are delivered via a variety of virtual and physical workloads across the entire enterprises. Besides that, the APPs are built and updated continuously and rapidly to meet the demands of business modes. Under the background of applications and microservices revolution, the network is expected to be a growing set of nerve clusters that can be anywhere along the cloud-edge-terminals. The network should understand how innovative application models work and dynamically extend the applications to wherever the applications are hosted.
CISILION summarizes four primary objectives for the new network in terms of aligning to the business, abstracting complexity, assuring performance and reducing risks [44]. The network in IIoT intelligence is expected to achieve some similar characteristics. As presented in Fig. 5, which refers from [46] and Prediction from TF security [47], the network connecting human-machine-things helps overcome the serious problem of information island, provides high-speed, wide-converage and low-latency communication, and supports the development of overlay applications, such as industrial big-data mining and immediate decisions. To be specific, in the increasingly demanding manufacturing environment, the network needs to be able to adapt quickly to changing manufacturing requirements. The network should provide convenient access of fast-changing set of equipments, staff, applications and services, so that resource scheduling and management can be focus more on. Providing steady continuous service without interruption is also essential for the network used in the manufacturing.
@@ IV-C2 Connotation
The industrial software is the product of the integration of industrial knowledge and information technologies. The product is updating iteratively as the long-term accumulation and application of industrial innovative knowledge. According to the functions and characteristics, the industrial software can be divided into four categories: R&D [^3] Research and Development software, production control software, information management software and embedded software[56]. Fig. 6 presents the primary functions and application instances of the four types of industrial softwares.
The industrial softwares help managers enhance control for the whole industrial processes and optimize production and operation schedules. For example, INTMES [^4] http://www.intmes.com/, whose primary business is designing personalized software solution for companies, provides a customized MES/SCM software555details shown in http://www.intmes.com/solution-content-CASE-2.html to Hanteng automobile Co.Ltd, a vehicle manufacturing company. With the help of MES/SCM, Hanteng’s workshop supports transparent management and fine control, and its labour efficiency is improved as well.
@@
The industrial software is highly complex for some reasons. First of all, there are various industries, among which process industry and discrete industry contain 39 major categories, 191 medium-granularity categories and 525 fine-granularity categories. Secondly, many professional knowledge and experience in machinery, electronics, optics, acoustics, sound control, fluid heat treatment are required to support research and development of industrial softwares. Thirdly, the complexity of industrial products is wide-span, from simple products such as clothing and toys to complex products in aviation and satellite. Moreover, the industrial softwares engage in all stages of industrial processes, while the industrial process is complex, including R&D, production, marketing, operation, maintenance, supply chain management and others. Deployed equipments are also various, functions involve production, implementation, test and measurement. Then, it is difficult for manufacturers to cooperate. For example, Boeing747 is jointly produced by more than 16000 enterprises with different scales from six countries. Finally, industrial software has high requirements for real-time reliability, because unreliable data is possible to cause huge problems in terms of cost and safety.
@@
In recent years, governments are promoting the rapid development of industrial software by policy incentives, at the same time, innovative technologies also provides opportunities to update industrial software. As the industry environment becomes more and more competitive and complex, the industrial software are upgrading in three ways, including systematic platform, cloud-based deployment and lightweight development. To be specific, early-stage industrial software are oriented specific scenarios, such as CAD for designing, CAE for simulation, ERP for enterprise management, etc. Current industry requires collaboration among multiple companies. Therefore, some systematic platforms are expected to support the cross-company interaction. The systematic platforms should integrate and utilize resources and tools to provide solutions for smart manufacturing. On the other hand, developing cloud-based industrial softwares [57][58] is an effective approach to deal with current operation models (IaaS, SaaS, PaaS) and customization demands. Guaranteeing data interconnection, real-time interaction and remote monitoring [21] is important topic in smart manufacturing, which promotes to develop lightweight mobile industrial APP.
@@ IV-D2 Connotation
Modeling methods that can build connections between physical world and cyber world f or IIoT belong to the modeling layer.
The digital twin is the most promising technology to realize the interactive between cyber spaces and physical spaces. The representation model of digital twin is shown in Fig. 7. The digital twin has five necessary keypoints: data, connection, model, interaction, and application. Consequently, the digital twin can run through the entire life cycle of physical processes, and provide capabilities with description, analysis, diagnosis, prediction, prescription, cognition, and so on [59][60]. Currently, the industrial-research-university communities explore the digital twin in two aspects, modeling and simulation. Modeling refers to establishing 3D virtual twins for physical entities. The twins are expected to reproduce appearance, geometric association, motion dynamics, and other attributes of the physical entities. The simulation is an activity to calculate, analyze and predict the future state of physical entities based on the virtual twins with industrial knowledge, mechanism, and data. As far as researches contributions at this stage, the modeling and simulation of digital twin depend on industrial software, including CAD, CAE, etc.
@@ IV-E2 Connotation
Manufacturing process planning determines how a product will be manufactured by selecting and sequencing manufacturing operations. The process planning aims at improving some specific performance (e.g., shorten processing time, minimize production cost) while satisfying a set of domain constraints [63] [64]. The manufacturing scheduling is the process of assigning manufacturing resources over time to the set of manufacturing processes in the process plan [63]. It is an optimization problem that schedules of resource allocation will be proposed with respect to criteria such as delay, throughput, or cost. The process planning and scheduling problem is becoming increasingly important to improve the flexible manufacturing. However, the problem is typically NP-hard, that is, it is impossible to get optimal solutions without the use of an essentially enumerative algorithm, while the computing complexity grows significantly as the problem scale increases.
@@ V A close look on lighthouse factories with IIoT intelligence
Since 2016, the World Economic Forum, in collaboration with McKinsey & Company, has been monitoring progress in advanced manufacturing worldwide. In 2017, the World Economic Forum identified 40 initial advanced manufacturing use cases that began to pilot intelligent technologies with some success. In 2018, 16 companies were recognized as leaders that achieve step-change results, both operational and financial, across individual sites [28]. In 2019, 28 additional facilities were identified as end-to-end lighthouses [29].
@@
Besides the successful end-to-end digital transformation lighthouse factories, some other collaborations are gradually released. Some important news are declared. Reports of ”How 5G Will Revolutionize Manufacturing” and ”Powering industries and a smart world with 5G, Cloud & AI” are published[111][112]. In 2019, Haier, China Mobile, Huawei and Mstar released the world’s first 5G smart factory joint solution [6]. The factory implements intelligence in terms of quality detection, maintenance, machine collaboration, material transportation, enery management and campus security under help of IIoT intelligence. In 2021, Midea Group, China Unicom, and Huawei rolled out Guangdong’s first 5G fully-connected smart manufacturing demonstration site at the factory of Midea’s Kitchen Appliance Division in Shunde district, Foshan city [7]. In the project, IIoT intelligence will enable achievements in multiple aspects. Specifically, the high-performance 5G dedicated network supports highly efficient production on the production line; the AI brain monitors and predicts robots’ behaviors and ensures that they are unified; machine vision is adopted to facilitate product quality inspection, significantly reducing the need for manual inspection; and the AGV carts deliver materials to specified locations.
@@ VI-A Industrial robots as intelligent labors
The industrial robot usually refers to the manufacturing-oriented multi-joint mechanical arms, or other mechanical devices with multiple axes. The primary characteristics involves automatically control, programmable, multi-purpose, fixed in place or mobile, etc. Based on the mechanical structure, the robots can be classified into six types, including articulated robot, SCARA robot, cartesian robot, Delta robot, cylindrical robot and others [113][23]. This paper presents some widely used robots. The articulated robot refers to a robot whose arm has at least three rotary joints. It has a large operating range, supports flexible movements and can grab objects close to the fuselage. The cartesian robot is a robot whose arm has three prismatic joints and whose axes are correlated with a cartesian coordinate system. The cartesian robots have simple kinematics, high positioning accuracy, and great stability. But its flexibility is poor, and the operating range is small. The Delta robot is a robot whose arms have concurrent prismatic or rotary joints. Its motor is mounted on a fixed base. By this way, the inertia of the robot in the process of motion can be reduced.
The industrial robots are mainly used to replace human in automatic and/or flexible production to some extent [22][23]. According to the statistics of operational sotck of industrial robots by customer industry and application - world made by IFR(the International Federation of Robotics)[36], the largest application of industries of industrial robots is automotive manufacturing, following that electronics is important as well. As a consequence, the automatic production and flexible manufacturing of these two areas are developing rapidly [31][28]. In terms of usage, most industrial robots are used for handling, followed by welding and assembling. These tasks usually require lots of physical energy and part of content has bad impacts on health.  At the same time, the statistics about annual installations also demonstrate that industrial robots have played a significant positive role in promoting industrial processes, and there is still large room for further development.
The most three important components of industrial robots are reducer, servo motor and controller. The reducer is the kernel of industrial robots, the servo provides energy for robots, and the controller likes the brain of robots. The tree primary components occupy 35%, 20% and 15% respectively, accounting for 70% in total. To further improve performance of robots, researchers and developers should break through technical barriers in these three parts.
@@ VI-B Machine vision system as eyes of industries
The goal of a machine vision system [114][115] is to create simplified models of the real world based on sensed images. These models should be understandable to computers and can restore real information (geometric shape, color, texture, category, printed matter, etc.) from images. Compared with human eyes, machine vision has significant differences and benefits, as presented in Table II. Thanks to these advantages, machine vision systems have been applied in industry.
@@
Machine vision systems have great differences between various industrial tasks. For general practical applications, there are some shared common performance requirements in terms of precision, speed, environmental robustness, and easy to use [116]. Without accurate measurement, the machine vision cannot successfully guide the robotic hand to perform automatic assembly of smartphone or mechanical watch  [117, 118, 119]. In the high-speed assembly, the average time of template matching is less than several microseconds for electronics manufacturing [120, 121]. A brief overview about how machine vision augments manufacturing technologies is given as follows.
The manufacturing techniques, including machining (milling, grinding, polishing), welding, injection molding, etc., can be improved with an auxiliary machine vision system. As an efficient and intelligent machining of complex free-form surface, robotic grinding demonstrates flexibility, intelligence and cost efficiency under the guidance of machine vision [122][123]. The raw workpiece is located by matching the measured point clouds and corresponding theoretical model. Thus, the machining error and allowance distribution for free-form surfaces can be optimized. Due to the accuracy requirement of grinding, accurate point clouds scanning and accurate calibrating between the machine vision system (scanning system) and the robot system are key to achieving high performance robotic grinding.
Equipped with machine vision system, the intelligence robotic welding is drawing attention from industrial community due to the high quality of product [124]. Robotic seam tracking is challenging due to the strong noise images(reflection, strong arc light and the large spatter), and a convolution neural network is proposed to extract the feature points of weld seam image [125]. A machine vision approach is introduced to tracking the welding electrode tip for assessing welding arc stability and polarity in manual metal arc [126]. The machine vision system thus becomes essential for welding monitor, diagnosis, prediction, and control. How to obtain useful information from noisy sequential images and how to use such information to improve the welding skill are two open questions for intelligent robotic welding [124].
With the help of machine vision, the robotic assembly has a long history due to the efficiency [127][128]. With the increasing capability of machine vision, the smart assembly demonstrates flexibility as well as challenges [129]. One challenge of robotic assembly is that the positional error of machine vision must meet the minimal requirement of individual task. Otherwise, the automatic assemblies may fail [130][131]. One direct solution is to improve the positional accuracy of machine vision by micro-measurement, such as eye-in-hand visual system [118] or optical configure of micro-view [132]. To achieve successful assembly, a data fusion trend with multi-modal sensors is considered by several recent works [133][134]. Regarding the automatic assembly, the challenges of machine vision are the accuracy requirement, multi-modal sensor fusion, deformable object manipulation, and advanced human-robot interaction.
Defect detection is highly recommended to ensure product quality following every manufacture/assembly procedure. The machine vision system is popular in defect detection due to its efficiency and attractive cost. The geometrical surface defect inspection is often achieved by point clouds registration between the scanned points and its template [135][136]. While, the assessment for compliant parts is still an open challenge [137][138]. The assembly defect detection, called as assembly line quality control, is to check the assembly correctness [139][140].
The on-line visual feedback of machine vision measurement (position, shape, velocity, etc.) enables the machining, welding, injection molding techniques with improved manufacturing quality and speed. Regarding the manufacturing of mechanical parts, the challenges of machine vision are the high accuracy requirement, harsh imaging conditions (occlusion, unwanted light, cutting chips, and cutting fluids), and field measurement of deformable/fluid materials.
@@ VI-C Networking facilitates ubiquitous connectivity
Current network is the product of the deep integration of the new generation of information, communication and modern industrial technologies. It has been one of the magic weapons to realize digital transformation of manufacturing. This paper will briefly introduces some related progresses of typical network technologies and applications, such as 5G, time sensitive networking, and identity resolution.
@@
The 5G network can connect autonomous industrial machines, ground vehicles, robots and software in an URLLC wireless communication, aiming at assistance of real-time control and management of industrial machinery towards smart factories [52][53][141]. The authors in [141] study 5G-enabled industrial Internet of Things, considering using machine learning to handle complex tasks at industrial machinery and 5G networks management,configuration,and control. Moreover, the authors design transfer learning based systems for Industrial Internet of Things (IIoT), and discuss how transfer learning can overcome the demand for high-quality large data samples required to training neural network in IIoT. An intelligent and green resource allocation mechanism for the IIoT under 5G heterogeneous networks is proposed [142], so as to ensure the high reliability and energy efficiency of IIoT services. Networking slicing is an essential technique to support massive connectivity and service isolation for 5G, so as to realize scale-up applications in the industry [54][55].
@@
Time-Sensitive Networking (TSN) [143] is a standardized technology that provides deterministic information transmission on traditional Ethernet. In [144], the authors introduce Open Platform Communication Unified Architecture TSN (OPC UA TSN) as a new technology and present the current view. In [145], the authors present an approach to integrating heterogeneous Industrial Ethernet applications over TSN, and deploy different real-time schedulers that are embedded in a Software-defined Network (SDN)[48][49] controller to allow different applications with various traffic patterns and Quality-of-Service (QoS) requirements to share the same network. In [146], the authors introduce the latest research progress of industrial communication systems, and analyze the technical challenges and future trends for TSN.
The identity resolution as an important network technology [147] builds the nerve endings of networking while used in industry. The kernel functions of the identity resolution are distribution, registration, management, analysis and routing of the identity of goods. Meanwhile, identity resolution can support the whole life cycle management of equipment, people and materials. It not only helps realize data inter-operability and data mining, but also provides a necessary condition for the intelligent management of manufacturing. There are many contributions about identity resolution systems [148, 149, 150]. Among which, RFID (radio frequency identification) based methods have made great progresses. RFID is an automatic object identification and data collection technology by using radio waves. Detail development of RFID from tags to middleware and application is discussed [151]. An insight to the application of RFID technology in the management of manufacturing and supply chain is summarized [152]. Some challenges in implementing the technology and potential future manufacturing application are discussed in the work as well [152]. With consideration of the information accuracy, time consuming of capturing and collection of real-time field information in the modern manufacturing, an affordable approach is introduced [150], which deploys wireless manufacturing technologies (such as RFID) for managing work-in-progress (WIP) inventories in manufacturing job shops.

### 22 Feb 2024 v2

@@ IV-B1 Motivations
The revolution of IIoT networks is inspired by global trends[53]. Firstly, because of more and more smart equipment and staff accessing the network in the domain of smart manufacturing, and the connection between systems, human, processes, locations and devices getting more distributed and complex, the economic value of the network for enterprises are growing. Secondly, the inherent unpredictability of business that the network quickly corrects the transmission modes to adjust to the new services, processes and models. Thirdly, many manufacturing processes are time sensitive and mission critical, which requires the communication time of network to be more precise and predictable[54].
Technology improvements encourage IIoT network revolution as well. The applications and data of enterprises are leaving the premises and moving to the public/private clouds. Part of the monolithic applications in many cases are being modularized into microservices that are delivered via a variety of virtual and physical workloads across the entire enterprises. Under the background of applications and microservices revolution, the network is expected to be a growing set of nerve clusters that can be anywhere along the cloud-edge-terminals.
CISILION summarizes four primary objectives for the new network in terms of aligning to the business, abstracting complexity, assuring performance and reducing risks [53]. The network in IIoT intelligence is expected to achieve some similar characteristics. As presented in Fig. 5, which refers from [55] and Prediction from TF security [56], the network connecting human-machine-things helps overcome the serious problem of information island, provides high-speed, wide-coverage and low-latency communication, and supports the development of overlay applications, such as industrial big-data mining and immediate decisions. To be specific, in the increasingly demanding manufacturing environment, the network needs to be able to adapt quickly to changing manufacturing requirements. The network should provide convenient access of fast-changing set of equipment, staff, applications and services, so that resource scheduling and management can be focus more on. Providing steady continuous service without interruption is also essential for the network used in the manufacturing.
@@ IV-C2 Connotation
The industrial software is the product of the integration of industrial knowledge and information technologies. The product is updating iteratively as the long-term accumulation and application of industrial innovative knowledge. According to the functions and characteristics, the industrial software can be divided into four categories: R&D software, production control software, information management software and embedded software[65]. Fig. 6 presents the primary functions and application instances of the four types of industrial softwares.
@@
The industrial software is highly complex for several reasons. First of all, there are various industries, among which process industry and discrete industry contain 39 major categories, 191 medium-granularity categories and 525 fine-granularity categories. Secondly, many professional knowledge and experience in machinery, electronics, optics, acoustics, sound control, fluid heat treatment are required to support research and development of industrial softwares. Thirdly, the complexity of industrial products is wide-span, from simple products such as clothing and toys to complex products in aviation and satellite. Moreover, the industrial softwares engage in all stages of industrial processes, while the industrial process is complex, including R&D, production, marketing, operation, maintenance, supply chain management and others. Deployed equipments are also various, functions involve production, implementation, test and measurement. Then, it is difficult for manufacturers to cooperate. For example, Boeing747 is jointly produced by more than 16000 enterprises with different scales from six countries. Finally, industrial software has high requirements for real-time reliability, because unreliable data is possible to cause huge problems in terms of cost and safety.
@@
In recent years, governments are promoting the rapid development of industrial software by policy incentives, at the same time, innovative technologies also provide opportunities to develop industrial software. As the industry environment becomes more and more competitive and complex, the industrial software is upgrading in three ways, including systematic platform, cloud-based deployment and lightweight development. To be specific, early-stage industrial software are oriented specific scenarios, such as CAD for designing, CAE for simulation, ERP for enterprise management, etc. Current industry requires collaboration among multiple companies. Therefore, some systematic platforms are expected to support the cross-company interaction. The systematic platforms should integrate and utilize resources and tools to provide solutions for smart manufacturing. On the other hand, developing cloud-based industrial softwares [66][67] is an effective approach to deal with current operation models (IaaS, SaaS, PaaS) and customization demands. Guaranteeing data interconnection, real-time interaction and remote monitoring [21] is important topic in smart manufacturing, which promotes to develop lightweight mobile industrial APP.
@@ IV-D2 Connotation
Modeling methods, which can build connections between physical world and cyber world for IIoT, belong to the modeling layer.
The digital twin is the most promising technology to realize the interactive between cyber spaces and physical spaces. Its representation model is shown in Fig. 7. The digital twin has five necessary keypoints: data, connection, model, interaction, and application. Consequently, the digital twin can run through the entire life cycle of physical processes, and provide capabilities with description, analysis, diagnosis, prediction, prescription, cognition, and so on [68][69]. Currently, the industrial-research-university communities explore the digital twin in two aspects, modeling and simulation. Modeling refers to establishing 3D virtual twins for physical entities. The twins are expected to reproduce appearance, geometric association, motion dynamics, and other attributes of the physical entities. The simulation is an activity to calculate, analyze and predict the future state of physical entities based on the virtual twins with industrial knowledge, mechanism, and data. As far as researches contributions at this stage, the modeling and simulation of digital twin depend on industrial software, including CAD, CAE, etc.
@@ IV-E2 Connotation
Manufacturing process planning determines how a product will be manufactured by selecting and sequencing manufacturing operations. The process planning aims at improving some specific performance (e.g., shorten processing time, minimize production cost) while satisfying a set of domain constraints [72, 73, 74]. The manufacturing scheduling is the process of assigning manufacturing resources over time to the set of manufacturing processes in the process plan [72]. It is an optimization problem that schedules of resource allocation will be proposed with respect to criteria such as delay, throughput, or cost. The process planning and scheduling problem is becoming increasingly important to improve the flexible manufacturing. However, the problem is typically NP-hard, that is, it is impossible to get optimal solutions without the use of an essentially enumerative algorithm, while the computing complexity grows significantly as the problem scale increases.
@@ V A close look on lighthouse factories with IIoT intelligence
Since 2016, the World Economic Forum, in collaboration with McKinsey & Company, has been monitoring progress in advanced manufacturing worldwide. In 2017, the World Economic Forum identified 40 initial advanced manufacturing use cases that began to pilot intelligent technologies with some success. In 2018, 16 companies were recognized as leaders that achieve step-change results, both operational and financial, across individual sites [28]. In 2019, 28 additional facilities were identified as end-to-end (E2E) lighthouses [29]. Until 2023, the number of E2E lighthouses increases to 54, and 17 are also sustainability lighthouses [31].
@@
Besides the successful end-to-end digital transformation lighthouse factories, more and more collaborations are being gradually released. For example, in 2019, Haier, China Mobile, Huawei and Mstar released the world’s first 5G smart factory joint solution [6]. The factory implements intelligence in terms of quality detection, maintenance, machine collaboration, material transportation, enery management and campus security under help of IIoT intelligence. In 2021, Midea Group, China Unicom, and Huawei rolled out Guangdong’s first 5G fully-connected smart manufacturing demonstration site at the factory of Midea’s Kitchen Appliance Division in Shunde district, Foshan city [7]. In 2022, Siemens and GENERA announced a collaboration [121], which involves the integration of Siemens technology in GENERA digital light processing systems, including operational technology, information technology, and automation. In addition, GENERA digital light processing solutions will be fully integrated into Siemens simulation and planning tools for factory design. In 2023, Siemens and Intel released a collaboration [122] to advance semiconductor manufacturing production efficiency and sustainability. Leveraging their respective portfolios of cutting-edge IoT solutions, along with Siemens automation solutions, the collaboration aimed to enhance semiconductor manufacturing efficiency and sustainability across scopes 1-3 of the value chain.
@@ VI-A Industrial robots as intelligent labors
The industrial robot usually refers to the manufacturing-oriented multi-joint mechanical arms, or other mechanical devices with multiple axes. The primary characteristics involves automatically control, programmable, multi-purpose, fixed in place or mobile, etc. Based on the mechanical structure, the robots can be classified into six types, including articulated robot, SCARA robot, cartesian robot, Delta robot, cylindrical robot and others [123][23]. The industrial robots are mainly used to replace human in automatic and/or flexible production to some extent [22][23]. According to the statistics of operational sotck of industrial robots by customer industry and application - world made by IFR(the International Federation of Robotics)[45], the largest application of industries of industrial robots is automotive manufacturing, following that electronics is important as well. As a consequence, the automatic production and flexible manufacturing of these two areas are developing rapidly [39][28]. In terms of usage, most industrial robots are used for handling, followed by welding and assembling. These tasks usually require lots of physical energy and part of content has bad impacts on health.  At the same time, the statistics about annual installations also demonstrate that industrial robots have played a significant positive role in promoting industrial processes, and there is still large room for further development.
@@
In practical manufacturing processes, robotics has played a crucial role in improving production efficiency. For example, in the World Economic Forum’s global lighthouse network 2022 [30], Haier refrigerator factory in Qingdao, China, uses collaborative robots for flexible manufacturing, resulting in a 52% improvement in assembly efficiency. Similarly, Sany Heavy Industry in Changsha, China, employs robotics in logistics execution, leading to an 11% increase in the on-time delivery rate.
@@ VI-B Machine vision system as eyes of industries
The goal of a machine vision system [124, 125, 126] is to create simplified models of the real world based on sensed images. These models should be understandable to computers and can restore real information (geometric shape, color, texture, category, printed matter, etc.) from images. Compared with human eyes, machine vision has significant differences and benefits, as presented in Table II. Thanks to these advantages, machine vision systems have been applied in industry.
@@
Machine vision systems have great differences between various industrial tasks. For general practical applications, there are some shared common performance requirements in terms of precision, speed, environmental robustness, and easy to use [127]. Without accurate measurement, the machine vision cannot successfully guide the robotic hand to perform automatic assembly of smartphone or mechanical watch  [128, 129, 130]. In the high-speed assembly, the average time of template matching is less than several microseconds for electronics manufacturing [131, 132]. The manufacturing techniques, including machining (milling, grinding, polishing), welding, injection molding, etc., can be improved with an auxiliary machine vision system. Rrobotic grinding demonstrates flexibility, intelligence and cost efficiency under the guidance of machine vision [133][134]. Equipped with machine vision system, the performance of intelligence robotic welding improves a lot [135]. Robotic seam tracking is challenging due to the strong noise images(reflection, strong arc light and the large spatter), and a convolution neural network is proposed to extract the feature points of weld seam image [136]. A machine vision approach is introduced to tracking the welding electrode tips for assessing welding arc stability and polarity in manual metal arc [137].
In general, the on-line visual feedback of machine vision measurement, including position, shape, velocity, etc., enables the machining, welding, injection molding techniques with improved manufacturing quality and speed. For example, Agilent’s machine vision toolkit has enabled speed and scale through democratized access to advanced vision capabilities [31]. Unilever complemented machine vision supervision platform for people safety and food safety compliance, decreasing 78% unsafe behaviours [30].
@@ VI-C Networking facilitates ubiquitous connectivity
Current network is the product of the deep integration of the new generation of information, communication and modern industrial technologies. It has been one of the magic weapons to realize digital transformation of manufacturing. This paper briefly introduces some related progresses of typical network technologies and applications, such as 5G, time sensitive networking, and identity resolution.
@@
The 5G network can connect autonomous industrial machines, ground vehicles, robots and software in an URLLC wireless communication, aiming at assistance of real-time control and management of industrial machinery towards smart factories [61][62][138]. The authors in [138] study 5G-enabled industrial Internet of Things, considering using machine learning to handle complex tasks at industrial machinery and 5G networks management,configuration,and control. An intelligent and green resource allocation mechanism for the IIoT under 5G heterogeneous networks is proposed [139], so as to ensure the high reliability and energy efficiency of IIoT services. Networking slicing is an essential technique to support massive connectivity and service isolation for 5G, so as to realize scale-up applications in the industry [63][64]. In practice, Exquisite Automotive Systems has joined forces with China Unicom Hebei and Huawei to create the world’s first flexible 5.5G production line in their Baoding factory, where the service interruption time per year reduces from 60 hours to 5.26 mins, and the annual output value increases \$112 million [140].
@@
Time-Sensitive Networking (TSN) [141] is a standardized technology that provides deterministic information transmission on traditional Ethernet. In [142], the authors introduce Open Platform Communication Unified Architecture TSN (OPC UA TSN) as a new technology and present the current view. In [143], the authors present an approach to integrating heterogeneous Industrial Ethernet applications over TSN, and deploy different real-time schedulers that are embedded in a Software-defined Network (SDN)[57][58] controller to allow different applications with various traffic patterns and Quality-of-Service (QoS) requirements to share the same network. In [144], the authors introduce the latest research progress of industrial communication systems, and analyze the technical challenges and future trends for TSN. In practice, Moxa, which is a leading provider of industrial networking, computing, and automation solutions, has successfully integrated TSN into the wafer manufacturing process of ELS System Technology, a leading Taiwanese lithography solution provider [145].